# Spectraformer: A Unified Random Feature Framework for Transformer

### v2

@@ A.1 Future work, limitations, and broader impact
Due to the significant number of experimented models, we are unable to perform hyper-parameter fine-tuning. Future work can benefit from extending the scope of the current research to other tasks, i.e., other non-textual LRA tasks (image classification, path-finder, and path-x) and short-context datasets/ tasks such as GLUE [38]. Transformer-based architecture can also benefit from Spectraformer. Past random feature Transformer works [12, 13] have experimented with conformer and vision Transformer with ViT. [39] provides a good foundation to extend our work. Spectraformer, although currently experimented in the Transformer setting, can also extend to non-parametric kernel classification. It could also benefit from ablation studies like the comparison variance in [14] or in [40]
Given the complex nature of language models in the ethical, energy, and societal domains, our work should be used responsibly  [41]
@@ A.2.1 Alternative Kernelized Attentions
Nadaraya-Watson kernel estimator in integration form [42]: Given the Gaussian kernel, where the probability of a single variable and the joint probability of two variables are estimated using kernel density estimation with the Gaussian kernel, then we have the following equation:
@@
Equation 12 lays the foundation for FourierFormer - a Transformer with attention based on this formulation [42].
Gaussian decomposition [43]: the attention mechanism can also be decomposed into the Gaussian kernel directly from Equation 4 without appealing to the use of kernel estimator [43].
$A_{i}=\sum_{j=1}^{N}\frac{1}{Z_{l}(\boldsymbol{q}_{i},\boldsymbol{K})}exp(\frac{-||\boldsymbol{q}_{i}-\boldsymbol{k}_{j}||^{2}_{2}}{2\sqrt{d_{k}}})exp(\frac{||\boldsymbol{q}_{i}||^{2}_{p=2}+||\boldsymbol{k}_{j}||^{2}_{p=2}}{2\sqrt{d_{k}}})\boldsymbol{v}_{j}$ (13)
Equation 13 leads to Implicit Kernel Attention (IKA) - a Transformer with attention based on this formulation [43].
Non-local operation [44] is popular decomposition of attention in computer vision. This is the basis for Vision Transformer and related work. Non-local operation itself a generalisation of the non-local means used in image denoising [45]. Non-local operation is essentially an operation that operates on the entire input range, instead of a specific window which is characteristic of convolutional operations (see Equation 14). Indeed, the attention vector output $\boldsymbol{y}$ can be defined using $f$ which is the softmax (see Equation 15) in the non-local operation.
$\mathbf{\boldsymbol{y}}_{i}=\frac{1}{C(\boldsymbol{x})}\sum_{\forall j}f(\boldsymbol{x}_{i},\boldsymbol{x}_{j})g(\boldsymbol{x}_{j})$ (14)
$\mathbf{\boldsymbol{y}}=softmax(\boldsymbol{x}^{T}\boldsymbol{W}_{\theta}^{T}\boldsymbol{W}_{\theta}\boldsymbol{x})g(\boldsymbol{x})$ (15)
We will show that the non-local operation in Equation 14 is a generalisation of the Nadaraya-Watson kernel estimator. Specifically, when defining $C(\boldsymbol{x})=\sum_{l}\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{l}),f=\mathcal{K}(\boldsymbol{q}_{i},\boldsymbol{k}_{l}),g=\boldsymbol{v}_{j}$, then we obtain the original Nadaraya-Watson kernel estimator in Equation 4. The variety of kernel decompositions of attention demonstrate the robustness of the kernel interpretation of attention and hence opening up an alley in attention improvement through the lens of kernel.
@@ A.2.2 Kernel Approximation Beyond Random Features
Kernel approximation is presented in Section 2.3 through the use of random features. However, there has been other kernel approximation techniques outside of random features. These include greedy basis selection [46], divide and conquer [47, 48], and nyström methods [49] (which led to the development of Skyformer [16] and Nyströmformer [17]).
Figure 4: Spectraformer weight matrices. Red tick denotes techniques which have been explored in the literature. Green tick denotes techniques which are explored in this work.
@@ A.3 Technical Details of Weight Approximators
Base [19] is the direct sampling from a distribution $\boldsymbol{P}_{ij}\sim p$, corresponding to the kernel $\mathcal{K}$ without any additional adjustment, and in approximating the RBF, $P=G$ where $G$ is the normal distribution.
$W_{Base}=\frac{1}{\sigma}\boldsymbol{P}$ (16)
FastFoodF [28] makes use of Hadamard and diagonal matrices to speed up Gaussian matrices $G$ constructions for RFF in $O(n\log{d})$ (100 times faster) with $O(n)$ space (1000 times less space) given $n\geq d$ against Base. However, FastFoodF increases variance and approximation error and decreases the concentration bound.
$W_{FastFood_{L}}=\frac{1}{\sigma}\boldsymbol{SHG\Pi HB}$ (17)
@@
SCRF (Signed Circulant Matrix Projection) [50] maintains a high matrix construction speed and $O(s)$ space ($s=t*d$) whilst maintaining the variance of Base, an improvement over FastFoodF. It does this by making use of circulant matrices.
$\boldsymbol{W}_{SCRF}=[\boldsymbol{P}^{(1)},\boldsymbol{P}^{(2)},...,\boldsymbol{P}^{(t)}],t=d|D$ (18)
@@
where $\boldsymbol{P}^{(i)}$ is a signed circulant Gaussian matrix, with the first column of circulant random matrix $\boldsymbol{C}^{(i)_{[d]}}$ of $\boldsymbol{P}^{(i)}$ drawn randomly from a Gaussian distribution $\mathcal{N}(\boldsymbol{0},\boldsymbol{I}/\sigma^{2})$. The circulant matrix via Discrete Fourier Transform can be defined as:
$\begin{array}[]{cc}\boldsymbol{C}=\frac{1}{m}\boldsymbol{F}^{*}diag(\boldsymbol{Fc})\boldsymbol{F}\\ \boldsymbol{F}=[e^{i\frac{2\pi}{m}kn}]^{m-1}_{k,n=0}\end{array}$ (19)
@@
P-Model [51] generalizes over both circulant matrices (including SCRF), FastFoodF, along with other Gaussian and semi-Gaussian structured matrix methods.
$\boldsymbol{W}_{P_{model}}=[\boldsymbol{g^{T}P_{1}},\boldsymbol{g^{T}P_{2}},...,\boldsymbol{g^{T}P_{s}}]^{T}$ (20)
@@
where $\boldsymbol{g}\in\mathbb{R}^{a}$ is a Gaussian vector and a sequence of $m$ $\boldsymbol{P}_{i}\in\mathbb{R}^{a*d}$ matrices with L2 norm columns, where $P_{i}$ represents a P-model. Semi-Gaussian structured matrices including Toeplitz, Hankel, and other Toeplitz-like matrices are an open area of research for random matrices.
@@
ORF (Orthogonal RF) [9] decreases the approximation error significantly compared to Base. This is done via replacing $G$ with a properly scaled random orthogonal matrix. However, generating orthogonal matrices become costly quickly as the number of dimensions increases.
$\boldsymbol{W}_{ORF}=\frac{1}{\sigma}\boldsymbol{SQ}$ (21)
@@
where $\boldsymbol{Q}$ is a uniformly distributed random orthogonal matrix (on the Stiefel manifold) obtained from the QR decomposition of $\boldsymbol{G}$, the set of rows of $\boldsymbol{Q}$ forming a basis in $\mathbb{R}^{d}$, and $\boldsymbol{S}$ is a diagonal matrix with entries sampled i.i.d from the $\chi$-distribution with $d$ degrees of freedom, thus making the rows of $\boldsymbol{SQ}$ and $\boldsymbol{G}$ identically distributed.
@@
SORF (Structured ORF) [9] decreases time and space complexity of ORF (from $O(d^{2})$ to $O(d\log{d})$ with almost no extra memory cost) by imposing structure on the orthogonal matrices, inspired by structural methods. SORF is unbiased with large $d$. We replace $\boldsymbol{S}$ with $\sqrt{d}$ and $\boldsymbol{Q}$ with structured matrix $\boldsymbol{HD}_{1}\boldsymbol{HD}_{2}\boldsymbol{HD}_{3}$
$\boldsymbol{W}_{SORF}=\frac{\sqrt{d}}{\sigma}\boldsymbol{HD}_{1}\boldsymbol{HD}_{2}\boldsymbol{HD}_{3}$ (22)
@@
where $\boldsymbol{D}_{i}\in\mathbb{R}^{d*d},i=1,2,3$ is diagonal ’sign-flipping‘ matrices with diagonal entries sampled from the Rademacher distribution, $\boldsymbol{H}$ is the normalized Walsh-Hadamard matrix.
ROM (Random Orthogonal Embeddings) [52] generalizes SORF ($t=3$) to:
$\boldsymbol{W}_{ROM}=\frac{\sqrt{d}}{\sigma}\prod_{i=1}^{t}\boldsymbol{SD}_{i}$ (23)
@@
where $\boldsymbol{S}$ is a class of $L_{2}$-normalized versions of Kronecker product matrices, of which the Hadamard $\boldsymbol{H}$ is a representative of, and $\boldsymbol{D}$ being independent diagonal matrices. When $\boldsymbol{S}=\boldsymbol{H}$ and $t=3$, we obtain SORF used in our experiments. Another ROM variant is S-Hybrid, which we do not experiment with, given as
$\boldsymbol{W}_{S_{Hybrid}}=\boldsymbol{SD}_{k}^{(\mathcal{U})}\prod_{i=1}^{k-1}\boldsymbol{SD}_{i}^{(\mathcal{R})}$ (24)
@@
where $\boldsymbol{D}_{k}^{(\mathcal{U})}$ is a diagonal matrix with i.i.d. $Unif(\mathcal{S}^{1})$ on the unit circle of $\mathbb{C}$. When being used in the component function, instead of calculating $\phi(\boldsymbol{x})^{T}\phi(\boldsymbol{y})$ in Equation 10. We model the equation as $Real(\overline{\phi(\boldsymbol{x})}^{T}\phi(\boldsymbol{y}))$.
@@
SimRF (Simplex RF) [11] is developed to be a positive RF with the most optimal solution, the MSE of which is the lowest in the geometric methods. This is an alternative to the ROM-based geometric methods. Two variants of SimRF are: weight-independent and weight-dependent. SimRF is defined as:
$\boldsymbol{W}_{simp}=\boldsymbol{DSR}$ (25)
@@
where $\boldsymbol{D}\in\mathbb{R}^{d*d}=diag(\boldsymbol{\omega}_{i}),\boldsymbol{R}\in\mathbb{R}^{d*d}$ being a random orthogonal matrix drawn from Haar measure on $O(d)$ and $\boldsymbol{S}\in\mathbb{R}^{d*d}$ with row $\boldsymbol{s}_{i}$:
$\boldsymbol{s}_{i}=\begin{cases}\sqrt{\frac{d}{d-1}}\boldsymbol{e}_{i}-\frac{\sqrt{d}+1}{(d-1)^{3/2}}(1,...,1,0)^{T}&1\leq i<d\\ \frac{1}{\sqrt{d-1}}(1,1,...,1,0)^{T}&i=d\end{cases}$ (26)
@@
being unit vectors which are manifestly normalized and subtend obtuse angles. SimRF variants are defined based on how $\boldsymbol{\omega}_{i}$ is constructed. For SimRFindep, $\boldsymbol{\omega}_{i}\sim\chi_{d}$. For Simdep, we permit the random vector direction $\{\hat{\boldsymbol{\omega}}_{i}\}$ to be correlated with norms $\{\boldsymbol{\omega}_{i}\}$:
$\boldsymbol{\omega}_{i}=-\frac{\sum_{j\neq i}\omega_{j}}{||\sum_{j\neq i}\omega_{j}||_{2}}\boldsymbol{\omega}_{i},i=1,...,d$ (27)
@@
QMC (Quasi-Monte Carlo) [25] evaluates on a low discrepancy sequence (e.g., Halton, Sobol’ Faure, and Niederreiter) of points instead of random points in Monte Carlo. Although the approximation error is only reduced minimally, QMC has been shown to perform better than MC in high dimensions and does not have undesirable clustering effect. To calculate QMC, we first assume that $p(\boldsymbol{x})=\prod^{d}_{j=1}p_{j}(\boldsymbol{x}_{j})$ factorizes with respect to the dimensions with $p_{j}(.)$ being a univariate density function. Then we define:
$\Phi^{-1}(\boldsymbol{t})=\left(\Phi_{1}^{-1}\left(\boldsymbol{t}_{1}\right),\ldots,\Phi_{d}^{-1}\left(\boldsymbol{t}_{d}\right)\right)\in\mathbb{R}^{d}$ (28)
@@
where $\Phi_{j}$ being the cumulative distribution function of $p_{j}$, $\boldsymbol{t}_{1},\boldsymbol{t}_{2},...,\boldsymbol{t}_{s}\in[0,1]^{d}$ being a low discrepancy sequence, and $\boldsymbol{\omega}_{i}=\Phi^{-1}(\boldsymbol{t}_{i})$. We can thus transform the integral on $\mathbb{R}^{d}$ in Equation 9 to an integral on the unit cube $[0,1]^{d}$ as
$\mathcal{K}(\boldsymbol{x}-\boldsymbol{x}^{\prime})=\int_{[0,1]^{d}}\exp\big{(}\mbox{i}(\boldsymbol{x}-\boldsymbol{x}^{\prime})^{\!\top}\Phi^{-1}(\boldsymbol{t})\big{)}\mbox{d}\boldsymbol{t}$
@@
$\boldsymbol{W}_{\text{QMC}}=[\Phi^{-1}(\boldsymbol{t}_{1}),\Phi^{-1}(\boldsymbol{t}_{2}),\ldots,\Phi^{-1}(\boldsymbol{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}\;.$
QMC can be further improved by a sub-grouped based rank-one lattice construction which improved complexity [53].
MM (Moment Matching) [26] [7] improves over QMC by removing the undesirable clustering effect and having the same approximation error with less features. This is done by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$:
$\boldsymbol{W}_{\text{MM}}=[\widetilde{\Phi}^{-1}(\boldsymbol{t}_{1}),\widetilde{\Phi}^{-1}(\boldsymbol{t}_{2}),\ldots,\widetilde{\Phi}^{-1}(\boldsymbol{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}$
where $\widetilde{\Phi}^{-1}(\boldsymbol{t}_{i})=\tilde{\boldsymbol{A}}^{-1}(\Phi^{-1}(\boldsymbol{t}_{i})-\tilde{\boldsymbol{\mu}})$ can be constructed using moment matching with sample mean $\tilde{\boldsymbol{\mu}}=\frac{1}{s}\sum_{i=1}^{s}\Phi^{-1}(\boldsymbol{t}_{i})$ and the square root of the sample covariance matrix $\tilde{\boldsymbol{A}}$ satisfying $\tilde{\boldsymbol{A}}\tilde{\boldsymbol{A}}^{\!\top}=\text{Cov}(\Phi^{-1}(\boldsymbol{t}_{i})-\tilde{\boldsymbol{\mu}})$.
SSF (Spherical Structured Features) [54] improves over QMC by including rotation-invariant kernel as well as in terms of complexity. Rotation-invariant property suggests the construction of feature maps using spherical equal weight approximation using Riesz s-energy on a d-dimensional sphere $\mathbb{S}^{d}:=\{\boldsymbol{x}\in\mathbb{R}^{d+1}|||\boldsymbol{x}||_{2}=1\}$. Specifically, we construct $\{\boldsymbol{v}_{i}\}^{s}_{i=1}$ asymptotically uniformly distributed on the sphere and obtain $\boldsymbol{V}:=[\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{s}]\in\mathbb{S}^{(d\times s}$, that $\Phi^{-1}(t)$ uses the one-dimensional QMC point, we have the weight matrix:
$\boldsymbol{W}_{\text{SSF}}=[\Phi^{-1}(t)\boldsymbol{v}_{1},\Phi^{-1}(t)\boldsymbol{v}_{2},\ldots,\Phi^{-1}(t)\boldsymbol{v}_{s}]^{\!\top}\in\mathbb{R}^{s\times(d+1)},$
SGQ (Sparse Grid Quadrature) [27] evolves from GQ. GQ (Gaussian Quadrature) [27] assumes that the kernel $k$ factorizes with respect to the dimensions and thus can be approximated by a one-dimensional Gaussian quadrature rule. However the total number of points $s$ scale exponentially with the dimensions. However GQ suffers form the curse of dimensionality. This is alleviated using Smolyak rule resulting in SGQ. Assuming the third-degree SGQ using symmetric univariate quadrature points $\left\{-\hat{p}_{1},0,\hat{p}_{1}\right\}$ with weights $(\hat{a}_{1},\hat{a}_{0},\hat{a}_{1})$, then we have:
$\boldsymbol{W}_{\text{SGQ}}\!=\![\boldsymbol{0}_{d},\hat{p}_{1}\boldsymbol{e}_{1},\ldots,\hat{p}_{1}\boldsymbol{e}_{d},-\hat{p}_{1}\boldsymbol{e}_{1},\ldots,-\hat{p}_{1}\boldsymbol{e}_{d}]^{\!\top}\!\in\!\mathbb{R}^{(2d+1)\times d}\!$

### v3

@@ A.1 Future work, limitations, and broader impact
Due to the significant number of experimented models, we are unable to perform hyper-parameter fine-tuning. Future work can benefit from extending the scope of the current research to other tasks, i.e., other non-textual LRA tasks (image classification, path-finder, and path-x) and short-context datasets/ tasks such as GLUE (Linzen et al., 2018). Transformer-based architecture can also benefit from Spectraformer. Past random feature Transformer works (Likhosherstov et al., 2022; 2023) have experimented with conformer and vision Transformer with ViT. Jeevan & Sethi (2022) provides a good foundation to extend our work. Spectraformer, although currently experimented in the Transformer setting, can also extend to non-parametric kernel classification. It could also benefit from ablation studies like the comparison variance in Chowdhury et al. (2022) or in Yao et al. (2023)
Given the complex nature of language models in the ethical, energy, and societal domains, our work should be used responsibly (Weidinger et al., 2021).s
@@ A.2.1 Alternative Kernelized Attentions
Nadaraya-Watson kernel estimator in integration form (Nguyen et al., 2022): Given the Gaussian kernel, where the probability of a single variable and the joint probability of two variables are estimated using kernel density estimation with the Gaussian kernel, then we have Equation 12, leading to the formulation of FourierFormer.
@@
Gaussian decomposition (Song et al., 2021): the attention mechanism can also be decomposed into the Gaussian kernel directly from Equation 4 without appealing to the use of kernel estimator (Song et al., 2021). Equation 13 leads to Implicit Kernel Attention (IKA).
$A_{i}=\sum_{j=1}^{N}\frac{1}{Z_{l}(\bm{q}_{i},\bm{K})}exp(\frac{-||\bm{q}_{i}-\bm{k}_{j}||^{2}_{2}}{2\sqrt{d_{k}}})exp(\frac{||\bm{q}_{i}||^{2}_{p=2}+||\bm{k}_{j}||^{2}_{p=2}}{2\sqrt{d_{k}}})\bm{v}_{j}$ (13)
Non-local operation (Wang et al., 2018) is popular decomposition of attention in computer vision. This is the basis for Vision Transformer and related work. Non-local operation itself a generalisation of the non-local means used in image denoising (Buades et al., 2005). Non-local operation is essentially an operation that operates on the entire input range, instead of a specific window which is characteristic of convolutional operations (see Equation 14). Indeed, the attention vector output $\bm{y}$ can be defined using $f$ which is the softmax (see Equation 15) in the non-local operation.
$\mathbf{\bm{y}}_{i}=\frac{1}{C(\bm{x})}\sum_{\forall j}f(\bm{x}_{i},\bm{x}_{j})g(\bm{x}_{j})$ (14)
$\mathbf{\bm{y}}=softmax(\bm{x}^{T}\bm{W}_{\theta}^{T}\bm{W}_{\theta}\bm{x})g(\bm{x})$ (15)
We will show that the non-local operation in Equation 14 is a generalisation of the Nadaraya-Watson kernel estimator. Specifically, when defining $C(\bm{x})=\sum_{l}\mathcal{K}(\bm{q}_{i},\bm{k}_{l}),f=\mathcal{K}(\bm{q}_{i},\bm{k}_{l}),g=\bm{v}_{j}$, then we obtain the original Nadaraya-Watson kernel estimator in Equation 4. The variety of kernel decompositions of attention demonstrate the robustness of the kernel interpretation of attention and hence opening up an alley in attention improvement through the lens of kernel.
@@ A.2.2 Kernel Approximation Beyond Random Features
Kernel approximation is presented in Section 2.3 through the use of random features. However, there has been other kernel approximation techniques outside of random features. These include greedy basis selection (Smola & Schökopf, 2000), divide and conquer (Hsieh et al., 2014; Zhang et al., 2013), and nyström methods (Williams & Seeger, 2000) (which led to the development of Skyformer (Chen et al., 2021) and Nyströmformer (Xiong et al., 2021)).
Figure 2: Spectraformer weight matrices. Red tick denotes techniques which have been explored in the literature. Green tick denotes techniques which are explored in this work. Please refer to Appendix A.3 for references.
inline]MAYBE: Table here like table 2 comparison of different kernel space and time complexities combined with table 3 comparison of convergence rates and required random features for kernel approximation error in (Liu et al., 2022)
@@ A.3 Technical Details of Weight Approximators
inline]update the equations so that we use the same dimension for the number of components and so on inline]Generative should be DGM, add Base to the graph
@@
Base (Rahimi & Recht, 2007) is the direct sampling from a distribution $\bm{P}_{ij}\sim p$, corresponding to the kernel $\mathcal{K}$ without any additional adjustment, and in approximating the RBF, $P=G$ where $G$ is the normal distribution.
$W_{Base}=\frac{1}{\sigma}\bm{P}$ (16)
FastFoodF (Le et al., 2013) makes use of Hadamard and diagonal matrices to speed up Gaussian matrices $G$ constructions for RFF in $O(n\log{d})$ (100 times faster) with $O(n)$ space (1000 times less space) given $n\geq d$ against Base. However, FastFoodF increases variance and approximation error and decreases the concentration bound.
$W_{FastFood_{L}}=\frac{1}{\sigma}\bm{SHG\Pi HB}$ (17)
@@
SCRF (Signed Circulant Matrix Projection) (Feng et al., 2015) maintains a high matrix construction speed and $O(s)$ space ($s=t*d$) whilst maintaining the variance of Base, an improvement over FastFoodF. It does this by making use of circulant matrices.
$\bm{W}_{SCRF}=[\bm{P}^{(1)},\bm{P}^{(2)},...,\bm{P}^{(t)}],t=d|D$ (18)
@@
where $\bm{P}^{(i)}$ is a signed circulant Gaussian matrix, with the first column of circulant random matrix $\bm{C}^{(i)_{[d]}}$ of $\bm{P}^{(i)}$ drawn randomly from a Gaussian distribution $\mathcal{N}(\bm{0},\bm{I}/\sigma^{2})$. The circulant matrix via Discrete Fourier Transform can be defined as:
$\begin{array}[]{cc}\bm{C}=\frac{1}{m}\bm{F}^{*}diag(\bm{Fc})\bm{F}\\ \bm{F}=[e^{i\frac{2\pi}{m}kn}]^{m-1}_{k,n=0}\end{array}$ (19)
@@
P-Model (Choromanski & Sindhwani, 2016) generalizes over both circulant matrices (including SCRF), FastFoodF, along with other Gaussian and semi-Gaussian structured matrix methods.
$\bm{W}_{P_{model}}=[\bm{g^{T}P_{1}},\bm{g^{T}P_{2}},...,\bm{g^{T}P_{s}}]^{T}$ (20)
@@
where $\bm{g}\in\mathbb{R}^{a}$ is a Gaussian vector and a sequence of $m$ $\bm{P}_{i}\in\mathbb{R}^{a*d}$ matrices with L2 norm columns, where $P_{i}$ represents a P-model. Semi-Gaussian structured matrices including Toeplitz, Hankel, and other Toeplitz-like matrices are an open area of research for random matrices.
@@
ORF (Orthogonal RF) (Yu et al., 2016) decreases the approximation error significantly compared to Base. This is done via replacing $G$ with a properly scaled random orthogonal matrix. However, generating orthogonal matrices become costly quickly as the number of dimensions increases.
$\bm{W}_{ORF}=\frac{1}{\sigma}\bm{SQ}$ (21)
@@
where $\bm{Q}$ is a uniformly distributed random orthogonal matrix (on the Stiefel manifold) obtained from the QR decomposition of $\bm{G}$, the set of rows of $\bm{Q}$ forming a basis in $\mathbb{R}^{d}$, and $\bm{S}$ is a diagonal matrix with entries sampled i.i.d from the $\chi$-distribution with $d$ degrees of freedom, thus making the rows of $\bm{SQ}$ and $\bm{G}$ identically distributed.
@@
SORF (Structured ORF) (Yu et al., 2016) decreases time and space complexity of ORF (from $O(d^{2})$ to $O(d\log{d})$ with almost no extra memory cost) by imposing structure on the orthogonal matrices, inspired by structural methods. SORF is unbiased with large $d$. We replace $\bm{S}$ with $\sqrt{d}$ and $\bm{Q}$ with structured matrix $\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$
$\bm{W}_{SORF}=\frac{\sqrt{d}}{\sigma}\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$ (22)
@@
where $\bm{D}_{i}\in\mathbb{R}^{d*d},i=1,2,3$ is diagonal ’sign-flipping‘ matrices with diagonal entries sampled from the Rademacher distribution, $\bm{H}$ is the normalized Walsh-Hadamard matrix.
ROM (Random Orthogonal Embeddings) (Choromanski et al., 2017) generalizes SORF ($t=3$) to:
$\bm{W}_{ROM}=\frac{\sqrt{d}}{\sigma}\prod_{i=1}^{t}\bm{SD}_{i}$ (23)
@@
where $\bm{S}$ is a class of $L_{2}$-normalized versions of Kronecker product matrices, of which the Hadamard $\bm{H}$ is a representative of, and $\bm{D}$ being independent diagonal matrices. When $\bm{S}=\bm{H}$ and $t=3$, we obtain SORF used in our experiments. Another ROM variant is S-Hybrid, which we do not experiment with, given as
$\bm{W}_{S_{Hybrid}}=\bm{SD}_{k}^{(\mathcal{U})}\prod_{i=1}^{k-1}\bm{SD}_{i}^{(\mathcal{R})}$ (24)
@@
where $\bm{D}_{k}^{(\mathcal{U})}$ is a diagonal matrix with i.i.d. $Unif(\mathcal{S}^{1})$ on the unit circle of $\mathbb{C}$. When being used in the component function, instead of calculating $\phi(\bm{x})^{T}\phi(\bm{y})$ in Equation 10. We model the equation as $Real(\overline{\phi(\bm{x})}^{T}\phi(\bm{y}))$.
@@
SimRF (Simplex RF) (Reid et al., 2023) is developed to be a positive RF with the most optimal solution, the MSE of which is the lowest in the geometric methods. This is an alternative to the ROM-based geometric methods. Two variants of SimRF are: weight-independent and weight-dependent. SimRF is defined as:
$\bm{W}_{simp}=\bm{DSR}$ (25)
@@
where $\bm{D}\in\mathbb{R}^{d*d}=diag(\bm{\omega}_{i}),\bm{R}\in\mathbb{R}^{d*d}$ being a random orthogonal matrix drawn from Haar measure on $O(d)$ and $\bm{S}\in\mathbb{R}^{d*d}$ with row $\bm{s}_{i}$:
$\bm{s}_{i}=\begin{cases}\sqrt{\frac{d}{d-1}}\bm{e}_{i}-\frac{\sqrt{d}+1}{(d-1)^{3/2}}(1,...,1,0)^{T}&1\leq i<d\\ \frac{1}{\sqrt{d-1}}(1,1,...,1,0)^{T}&i=d\end{cases}$ (26)
@@
being unit vectors which are manifestly normalized and subtend obtuse angles. SimRF variants are defined based on how $\bm{\omega}_{i}$ is constructed. For SimRFindep, $\bm{\omega}_{i}\sim\chi_{d}$. For Simdep, we permit the random vector direction $\{\hat{\bm{\omega}}_{i}\}$ to be correlated with norms $\{\bm{\omega}_{i}\}$:
$\bm{\omega}_{i}=-\frac{\sum_{j\neq i}\omega_{j}}{||\sum_{j\neq i}\omega_{j}||_{2}}\bm{\omega}_{i},i=1,...,d$ (27)
@@
QMC (Quasi-Monte Carlo) (Avron et al., 2016) evaluates on a low discrepancy sequence (e.g., Halton, Sobol’ Faure, and Niederreiter) of points instead of random points in Monte Carlo. Although the approximation error is only reduced minimally, QMC has been shown to perform better than MC in high dimensions and does not have undesirable clustering effect. To calculate QMC, we first assume that $p(\bm{x})=\prod^{d}_{j=1}p_{j}(\bm{x}_{j})$ factorizes with respect to the dimensions with $p_{j}(.)$ being a univariate density function. Then we define:
$\Phi^{-1}(\bm{t})=\left(\Phi_{1}^{-1}\left(\bm{t}_{1}\right),\ldots,\Phi_{d}^{-1}\left(\bm{t}_{d}\right)\right)\in\mathbb{R}^{d}$ (28)
@@
where $\Phi_{j}$ being the cumulative distribution function of $p_{j}$, $\bm{t}_{1},\bm{t}_{2},...,\bm{t}_{s}\in[0,1]^{d}$ being a low discrepancy sequence, and $\bm{\omega}_{i}=\Phi^{-1}(\bm{t}_{i})$. We can thus transform the integral on $\mathbb{R}^{d}$ in Equation 9 to an integral on the unit cube $[0,1]^{d}$ as
$\mathcal{K}(\bm{x}-\bm{x}^{\prime})=\int_{[0,1]^{d}}\exp\big{(}\mbox{i}(\bm{x}-\bm{x}^{\prime})^{\!\top}\Phi^{-1}(\bm{t})\big{)}\mbox{d}\bm{t}$
@@
$\bm{W}_{\text{QMC}}=[\Phi^{-1}(\bm{t}_{1}),\Phi^{-1}(\bm{t}_{2}),\ldots,\Phi^{-1}(\bm{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}\;.$
QMC can be further improved by a sub-grouped based rank-one lattice construction which improved complexity (Lyu et al., 2020).
MM (Moment Matching) (Shen et al., 2017; Liu et al., 2022) improves over QMC by removing the undesirable clustering effect and having the same approximation error with less features. This is done by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$:
$\bm{W}_{\text{MM}}=[\widetilde{\Phi}^{-1}(\bm{t}_{1}),\widetilde{\Phi}^{-1}(\bm{t}_{2}),\ldots,\widetilde{\Phi}^{-1}(\bm{t}_{s})]^{\!\top}\in\mathbb{R}^{s\times d}$
where $\widetilde{\Phi}^{-1}(\bm{t}_{i})=\tilde{\bm{A}}^{-1}(\Phi^{-1}(\bm{t}_{i})-\tilde{\bm{\mu}})$ can be constructed using moment matching with sample mean $\tilde{\bm{\mu}}=\frac{1}{s}\sum_{i=1}^{s}\Phi^{-1}(\bm{t}_{i})$ and the square root of the sample covariance matrix $\tilde{\bm{A}}$ satisfying $\tilde{\bm{A}}\tilde{\bm{A}}^{\!\top}=\text{Cov}(\Phi^{-1}(\bm{t}_{i})-\tilde{\bm{\mu}})$.
SSF (Spherical Structured Features) (Lyu, 2017) improves over QMC by including rotation-invariant kernel as well as in terms of complexity. Rotation-invariant property suggests the construction of feature maps using spherical equal weight approximation using Riesz s-energy on a d-dimensional sphere $\mathbb{S}^{d}:=\{\bm{x}\in\mathbb{R}^{d+1}|||\bm{x}||_{2}=1\}$. Specifically, we construct $\{\bm{v}_{i}\}^{s}_{i=1}$ asymptotically uniformly distributed on the sphere and obtain $\bm{V}:=[\bm{v}_{1},\bm{v}_{2},\ldots,\bm{v}_{s}]\in\mathbb{S}^{(d\times s}$, that $\Phi^{-1}(t)$ uses the one-dimensional QMC point, we have the weight matrix:
$\bm{W}_{\text{SSF}}=[\Phi^{-1}(t)\bm{v}_{1},\Phi^{-1}(t)\bm{v}_{2},\ldots,\Phi^{-1}(t)\bm{v}_{s}]^{\!\top}\in\mathbb{R}^{s\times(d+1)},$
SGQ (Sparse Grid Quadrature) (Dao et al., 2017) evolves from GQ. GQ (Gaussian Quadrature) (Dao et al., 2017) assumes that the kernel $k$ factorizes with respect to the dimensions and thus can be approximated by a one-dimensional Gaussian quadrature rule. However the total number of points $s$ scale exponentially with the dimensions. However GQ suffers form the curse of dimensionality. This is alleviated using Smolyak rule resulting in SGQ. Assuming the third-degree SGQ using symmetric univariate quadrature points $\left\{-\hat{p}_{1},0,\hat{p}_{1}\right\}$ with weights $(\hat{a}_{1},\hat{a}_{0},\hat{a}_{1})$, then we have:
$\bm{W}_{\text{SGQ}}\!=\![\bm{0}_{d},\hat{p}_{1}\bm{e}_{1},\ldots,\hat{p}_{1}\bm{e}_{d},-\hat{p}_{1}\bm{e}_{1},\ldots,-\hat{p}_{1}\bm{e}_{d}]^{\!\top}\!\in\!\mathbb{R}^{(2d+1)\times d}\!$